SAS Computers owns a patent on a computer processor. The processor was developed and capitalized at a cost of €2,030,000 in the beginning of 2015. It was expected to be economically useful for 7 years and have no residual value. At the beginning of 2018, a new processor was developed, making the old processor worth €850,000 (independent appraiser) with €50,000 total cost to sell. The present value of the processor’s future cash flows, given the development of the newer processor, is estimated to be €870,000. At this point, it is expected to have a useful life of 4 years with no residual value. Is the processor impaired in 2018? If it is impaired, prepare the journal entry to record the loss. Also prepare the journal entry for amortization in 2018. Show your work. How much is the patent worth at the end of 2018?
Answer #1

Ans:

Calculation of the book value of the processor at the beginning of the year 2018
Total cost of the processor (A) $ 2,030,000
Life of the processor (B) 7
Amortisation cost per annum (C=A/B) $ 290,000
Amortisation from 2015 to 2017 (3 years *$300,000) (D) $ 870,000
Balance book value of the processor (E=A-D) $ 1,160,000
Calculation of the value of the processor after development of new processor
Processor value at the beginning of the year 2018 $ 850,000
Less: Cost to sell $ 50,000
Balance value of the processor at the beginning of the year 2018 (F) $ 800,000
Present value of future cash flows (G) $ 870,000
Value of the processor at the beginning of the year 2018 (H= lower of F and G) $ 800,000
Ans Impairment loss = 1,160,000-800,000 (E-H) $ 360,000
Journal entries
Impairment loss:
Date Particulars Debit Credit
01-Jan Impairment:Loss $ 360,000
Processor $ 360,000
(To record the Impairment Loss)
Amortisation for the year 2018
31-Dec Amortisation expenses (800,000/4 years) $ 200,000
Processor $ 200,000
(To record the Current Year Amortisation Expenses)